# Logarithmic or algebraic: roughening of an active Kardar-Parisi-Zhang surface

### v3

@@ C.4 One-loop corrections to $\lambda,\lambda_{1}$ and $\lambda_{2}$
Fig. 6 shows representative one-loop Feynman diagrams for renormalisation of $\lambda$, $\lambda_{1}$ and $\lambda_{2}$ at the one-loop order, where $a,b,c$ can take values $0,1,2$, and $\lambda_{0}=\lambda$. There are many possible similar diagrams depending on $a,b,c$ but for a particular set of vertices (i.e for a fixed values of $a,b,c$) there exist two diagrams that may contribute to the corrections of any of $\lambda,\lambda_{1},\lambda_{2}$ as shown in Fig. 6. When ($a=b=c=1$) then Fig. 6$(a)$ has symmetry factor $24$ and contributes following,
$\frac{\lambda_{1}^{3}}{3!}\times 24\times\Bigr{[}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}\hat{h}\nabla_{m}h\nabla_{t}h\Bigr{]}\int_{{\bf q},\Omega}q_{i}q_{j}q_{m}q_{t}$
×$\quad\quad\quad\quad\quad\times\frac{2D}{(\Omega+i\nu q^{2})^{2}(\Omega-i\nu q^{2})^{2}}$ (46)
$=\frac{48\times\lambda_{1}^{3}}{3!\times 8}\times\Bigr{[}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}\hat{h}\nabla_{m}h\nabla_{t}h\Bigr{]}\frac{2D}{8\nu^{3}}\int\frac{d^{2}q}{q^{2}}$
×$\quad\quad\quad\quad\quad\times\Bigr{[}\delta_{ij}\delta_{mt}+\delta_{im}\delta_{jt}+\delta_{it}\delta_{jm}\Bigr{]}$ (47)
Similarly for ($a=b=c=1$) Fig. 6$(b)$ has symmetry factor $48$ and contributes following,
$\frac{48\lambda_{1}^{3}}{3!}\times\Bigr{[}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}\hat{h}\nabla_{m}h\nabla_{t}h\Bigr{]}\int_{{\bf q},\Omega}q_{i}q_{j}q_{m}q_{t}$
×$\quad\quad\quad\quad\quad\times\frac{2D}{(\Omega^{2}+\nu^{2}q^{4})(\Omega+i\nu q^{2})^{2}}$ (48)
$=-\frac{48\times\lambda_{1}^{3}}{3!\times 8}\times\Bigr{[}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}\hat{h}\nabla_{m}h\nabla_{t}h\Bigr{]}\frac{2D}{8\nu^{3}}\int\frac{d^{2}q}{q^{2}}$
×$\quad\quad\quad\quad\quad\times\Bigr{[}\delta_{ij}\delta_{mt}+\delta_{im}\delta_{jt}+\delta_{it}\delta_{jm}\Bigr{]}$ (49)
We can see that Eq. (49) is exactly same as Eq. (47) but with a negative sign. They cancel each other contributing nothing to the corrections of $\lambda,\lambda_{1}$ and $\lambda_{2}$. In fact, for any combination of $a,b$ and $c$ there exist diagrams like Fig. 6$(a)$ and Fig. 6$(b)$ added up to zero contributing nothing to $\lambda,\lambda_{1}$ and $\lambda_{2}$.. This shows that there are no relevant corrections to $\lambda,\lambda_{1}$ and $\lambda_{2}$ at the one-loop order.
C.5 RG flow equations and scaling
@@
$\nu^{<}=\nu\Bigr{[}1+\Bigl{(}\frac{\lambda_{1}^{2}D}{2\nu^{3}}+\frac{5}{8}\frac{\lambda\lambda_{1}D}{\nu^{3}}+\frac{\lambda_{2}^{2}D}{8\nu^{3}}\Bigl{)}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{d^{2}q}{q^{2}}\Bigr{]}$
. We set $b=e^{\delta l}\approx 1+\delta l$ and define dimensionless coupling constants by $g=\frac{\lambda^{2}D}{\nu^{3}}\frac{1}{2\pi}$,$\gamma_{1}=\frac{\lambda_{1}}{\lambda}$ and $\gamma_{2}=\frac{\lambda_{2}}{\lambda}$ to obtain the RG flow equations in the main text.
In dimensions greater than 2 (where $\gamma_{2}=0$), we set $d=2+\epsilon$ with $\epsilon>0$. We aim to determine the fixed point and the corresponding scaling exponents accurately up to $O(\epsilon)$ only. To achieve this, it suffices to substitute $d=2$ in the one-loop integrals. For example, in a general dimension where $d>2$, the RG flow equation governing the parameter $g$ is expressed as follows:
$\frac{dg}{dl}=$ $g\Biggr{[}2-d+g\biggl{\{}\frac{1}{4}+\frac{3}{d(d+2)}\gamma_{1}^{2}+\frac{\gamma_{1}}{d}-\frac{3}{4}\frac{2-d}{d}-\frac{3}{d}\gamma_{1}^{2}$
$-\frac{3\gamma_{1}}{d(d+2)}\biggl{(}3-\frac{d+2}{2}+\frac{d(d+2)}{2}\biggl{)}\biggl{\}}\Biggr{]}$
@@
$\frac{dg}{dl}=-\epsilon g-{\cal\tilde{A}}(\gamma_{1})g^{2}.$
Consequently, this produces $g$ correctly to ${\cal O}(\epsilon)$; see, e.g., Ref. Barabási and Stanley (1995).
Appendix D Correlation function

### v4

@@ III.4 One-loop corrections to $\lambda,\lambda_{1}$ and $\lambda_{2}$
Fig. 6 shows representative one-loop Feynman diagrams for renormalisation of $\lambda$, $\lambda_{1}$ and $\lambda_{2}$ at the one-loop order, where $a,b,c$ can take values $0,1,2$, and $\lambda_{0}=\lambda$. There are many possible similar diagrams depending on $a,b,c$ but for a particular set of vertices (i.e for a fixed values of $a,b,c$) there exist two diagrams that may contribute to the corrections of any of $\lambda,\lambda_{1},\lambda_{2}$ as shown in Fig. 6. When ($a=b=c=1$) then Fig. 6(a)(a) has symmetry factor $24$ and contributes following,
$\frac{\lambda_{1}^{3}}{3!}\times 24\times\Bigr[\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}\hat{h}\nabla_{m}h\nabla_{t}h\Bigr]\int_{{\bf q},\Omega}q_{i}q_{j}q_{m}q_{t}\times\frac{2D}{(\Omega+i\nu q^{2})^{2}(\Omega-i\nu q^{2})^{2}}$ (46)
$=\frac{48\times\lambda_{1}^{3}}{3!\times 8}\times\Bigr[\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}\hat{h}\nabla_{m}h\nabla_{t}h\Bigr]\frac{2D}{8\nu^{3}}\int\frac{d^{2}q}{q^{2}}\times\Bigr[\delta_{ij}\delta_{mt}+\delta_{im}\delta_{jt}+\delta_{it}\delta_{jm}\Bigr]$ (47)
Similarly for ($a=b=c=1$) Fig. 6(b)(b) has symmetry factor $48$ and contributes following,
$\frac{48\lambda_{1}^{3}}{3!}\times\Bigr[\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}\hat{h}\nabla_{m}h\nabla_{t}h\Bigr]\int_{{\bf q},\Omega}q_{i}q_{j}q_{m}q_{t}\times\frac{2D}{(\Omega^{2}+\nu^{2}q^{4})(\Omega+i\nu q^{2})^{2}}$ (48)
$=-\frac{48\times\lambda_{1}^{3}}{3!\times 8}\times\Bigr[\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}\hat{h}\nabla_{m}h\nabla_{t}h\Bigr]\frac{2D}{8\nu^{3}}\int\frac{d^{2}q}{q^{2}}\times\Bigr[\delta_{ij}\delta_{mt}+\delta_{im}\delta_{jt}+\delta_{it}\delta_{jm}\Bigr]$ (49)
We can see that Eq. (49) is exactly same as Eq. (47) but with a negative sign. They cancel each other contributing nothing to the corrections of $\lambda,\lambda_{1}$ and $\lambda_{2}$. In fact, for any combination of $a,b$ and $c$ there exist diagrams like Fig. 6(a)(a) and Fig. 6(b)(b) added up to zero contributing nothing to $\lambda,\lambda_{1}$ and $\lambda_{2}$.. This shows that there are no relevant corrections to $\lambda,\lambda_{1}$ and $\lambda_{2}$ at the one-loop order.
III.5 RG flow equations and scaling
@@
$\nu^{<}=\nu\Bigr[1+\Bigl(\frac{\lambda_{1}^{2}D}{2\nu^{3}}+\frac{5}{8}\frac{\lambda\lambda_{1}D}{\nu^{3}}+\frac{\lambda_{2}^{2}D}{8\nu^{3}}\Bigl)\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{d^{2}q}{q^{2}}\Bigr].$
We set $b=e^{\delta l}\approx 1+\delta l$ and define dimensionless coupling constants by $g=\frac{\lambda^{2}D}{\nu^{3}}\frac{1}{2\pi}$$\gamma_{1}=\frac{\lambda_{1}}{\lambda}$ and $\gamma_{2}=\frac{\lambda_{2}}{\lambda}$ to obtain the RG flow equations in the main text.
In dimensions greater than 2 (where $\gamma_{2}=0$), we set $d=2+\epsilon$ with $\epsilon>0$. We aim to determine the fixed point and the corresponding scaling exponents accurately up to $O(\epsilon)$ only. To achieve this, it suffices to substitute $d=2$ in the one-loop integrals. For example, in a general dimension $d>2$, the RG flow equation governing the parameter $g$ is expressed as follows:
$\frac{dg}{dl}=g\Biggr[2-d+g\biggl\{\frac{1}{4}+\frac{3}{d(d+2)}\gamma_{1}^{2}+\frac{\gamma_{1}}{d}-\frac{3}{4}\frac{2-d}{d}-\frac{3}{d}\gamma_{1}^{2}-\frac{3\gamma_{1}}{d(d+2)}\biggl(3-\frac{d+2}{2}+\frac{d(d+2)}{2}\biggl)\biggl\}\Biggr]$
@@
$\frac{dg}{dl}=-\epsilon g-{\cal\tilde{A}}(\gamma_{1})g^{2},$
that is correct to ${\cal O}(\epsilon)$. Consequently, this produces $g$ correctly to ${\cal O}(\epsilon)$; see, e.g., Ref. Barabási and Stanley (1995).
IV Correlation function